OpenStudy (anonymous):

Write a slope intercept eqation for the line that passes through (-3,-15) and is perpendicular to -6x + 8y = 3. Show all work.

OpenStudy (anonymous):

First change -6x + 8y = 3 to slope-intercept form. -6x + 8y = 3 8y = 6x + 3 y = 6/8x + 3 y = 3/4x + 3 The slope of a perpendicular line is the opposite recipricol: so 3/4 changes to -4/3. Now to make the new equation we use the point-slope form: y - y1 = m(x - x1 slope(m) = -4/3 x = -3 y = -15 So: y -(-15) = -4/3(x -(-15) y + 15 = -4/3(x + 15) y + 15 = -4/3x - 20 y = -4/3x - 35

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yes it did ty so much